OpenStudy (anonymous):

Why is (6, –2) not a solution to the system of equations given? 2x-3y= 18 5x=3y=3

OpenStudy (mikurout):

Is your 2nd equation correct? 5x = 3y = 3

OpenStudy (anonymous):

if it is, then it's because that second equation would simplify to y=1 and x=3/5 and x=/=y so it all doesnt work. but i suspect one of those should be subtraction

OpenStudy (anteater):

I bet it is 5x + 3y = 3, since the + and the = are on the same key.

OpenStudy (anonymous):

oooh good guess

OpenStudy (anonymous):

Why is (6, –2) not a solution to the system of equations given? 2x-3y= 18 5x+3y=3

OpenStudy (anteater):

So, if you were to graph the lines, what would the "solution" look like on the graph?

OpenStudy (anonymous):

alternately, plug in the values x=6 and y=-2 into the second equation and see if its a valid solution

OpenStudy (anteater):

Yes, the "solution" to a system of linear equations (or any other system) is the point (or points) that the graphs of the functions have in common. In other words, the points where they intersect.

OpenStudy (anteater):

And, as LifeEngineer pointed out, it can be verified algebraically by substituting the values for x and y into the equations to see if they "work" in both equations.

OpenStudy (anonymous):

Guys I'm calling it a night. good night and god bless

OpenStudy (anteater):

So 2x - 3y = 18 ---> 2(6) - 3(-2) = 12 + 6 = 18 (so (6,-2) works in that equation) and 5x + 3y = 3 ---> 5(6) + 3(-2) = 30 - 6 = 24 (so, since 3 isn't equal to 24, the point (6, -2) doesn't appear to work in that equation.)

OpenStudy (anteater):

Have a good night! :D

OpenStudy (anteater):

You too, LifeEngineer. :)

OpenStudy (anonymous):

Good night to all :D